OpenStudy (anonymous):

Solve -2(-3x - 8) = 4. -3 -2 2 3

OpenStudy (anonymous):

6x+16=4 6x=4-16 6x=-12 divide both sides by 6. so, x=-2. :)

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

u're welcome! :D